Company A buys 70% of 20000 2€ordinary shares of company B for 5€ per share.

The Noncurrent Assets of company B have a Fair value of 40000€.

Company A Company B
Noncurrent Assets 90000 20000
Investment -
Net Current Assets 18000 28000
Net Assets 150000 48000
Share capital 124000 40000
Retained Earnings 26000 8000
Tot Equity 150000 48000

Compute the amount of Goodwill and Non controlling interest
Answer #1

Fair value of Company B as on date of acquisition : Non Current Assets + Net Current Assets = 40,000 + 28000 = 68,000

So the amount of goodwill = Amount Paid - Fair value of assets acquired = (20000*70%*2) - (68000*20%) = 13600 - 70000 = 56,400

Non Controlling Interest = Fair value of the company B * Non controlling interest = 68000 * 30% = 20,400.